# GenNBV: Generalizable Next-Best-View Policy for Active 3D Reconstruction

### v2

@@ 1 Introduction
Recent advances in 3D reconstruction [26, 40, 41] and neural rendering [23, 39, 50] have significantly enhanced the quality of 3D digitization of large scenes, such as buildings and city landmarks [12, 49, 21, 45, 20]. However, image-capturing process still remains time-consuming and labor-intensive. To scan a building-scale scene using a drone, it may take several days’ effort of a professional team to ensure a full coverage. Moreover, even professional pilots may miss some areas at the first scan, leading to multiple rounds for rescanning.
To alleviate the effort in manual scanning, active 3D reconstruction algorithms, especially the Next-Best-View (NBV) policy, have emerged as a promising approach to automate view planning. However, typical rule-based NBV policies [17, 25, 48, 28, 10, 30] heavily rely on empirically designed view-selection criteria and hand-crafted action space, such as a hemisphere as shown in Fig. 1, which limit their generalizability to unseen scenes. To enhance generalization capacity, pioneering works on learning-based NBV policy have been proposed [28, 48, 5, 3]. Nevertheless, they are constrained into impractical setups in real-world scenarios, such as object-centric capturing and limited action space for ground robots, resulting in incomplete 3D reconstruction because of significant self-occlusion.
@@
To overcome these two challenges, we propose GenNBV, an end-to-end generalizable NBV policy. It has a novel design of action spaces and state embeddings, which allows free-space exploration and ensures a large coverage when applying to unseen objects. We firstly extend former limited action space, like a hemisphere, to 5D free space. As shown in Fig. 1, this new space is composed of a position subspace of approximately $20m$ x $20m$ x $10m$ and an omnidirectional heading subspace. This free space enables our agent drone to scan from any viewpoint. Moreover, with a much larger action space, our RL-based policy can be easily generalized to building-scale geometries without concerning the potential scale shift from training to evaluation. On the contrary, many hand-crafted designs [25, 17] limit available viewpoints to a small predefined set, and thus cannot generalize if there is a drastic change in scales.
Second, in order to generalize a well-trained NBV policy to unseen scenes during evaluation, we propose a novel state representation directly from raw sensory inputs to guide the policy. In contrast to the widely-used neural radiance field (NeRF) [25, 48, 17, 36, 30], our representation does not need time-consuming per-scene optimization. Specifically, our NBV policy incorporates a multi-source state embedding, which comprises three key components: a novel geometric embedding, a semantic embedding, and an action embedding. The geometric embedding is derived from multi-view depth maps and represents an encoded probabilistic 3D grid, while the semantic embedding is extracted from RGB images, and the action embedding is an encoded viewpoint sequence. In contrast to 2D representations used in previous RL-based NBV policies [5, 3], our multi-source state embedding, extracted from a probabilistic 3D grid, is a more comprehensive representation that supports a robust prediction of next viewpoints.
To validate the effectiveness of GenNBV, we construct a benchmark for training and evaluation with Houses3K [27] dataset from the NVIDIA Isaac Gym [22] simulator. We further test its generalization ability on novel buildings and object categories from OmniObject3D [44] and Objaverse [7] datasets. Our evaluation metrics quantify the completeness, accuracy, and efficiency for active 3D reconstruction. To consolidate completeness and efficiency into a single number, we also propose a metric, the Area Under the Curve (AUC) of the coverage. Compared with heuristic [13], information gain-based [13, 17, 48], and RL-based baselines [28], our method achieves the best result under different metrics. We also provide visualizations to demonstrate the generalizability of GenNBV. A demo video is attached in the appendix to further illustrate our method.
@@ 3.2 Generalizable State Embedding
To ensure that the policy learned on one set of 3D objects can generalize to objects with different appearances and structures, a smart state embedding of raw sensor observations is needed to capture invariant features across different objects. Previous methods  [28, 3] only extract representations in 2D space, which are very sensible to appearance changes, and our experiments have shown that policies only trained on 2D features may not generalize to different objects. Instead, we propose a multi-source representation that has better generalizability.
@@
During each scanning (one episode in RL), at each step $t+1$, we update the probabilistic occupancy grid $F_{t+1}^{G}$ based on the preceding grid $F_{t}^{G}$ and current observation. Specifically, the grid is updated through Bresenham’s line algorithm [2], which casts the ray path in 3D space between the camera viewpoint and the endpoints among the point cloud back-projected from depth $D_{t+1}$. Following the classical occupancy grid mapping algorithm [42], we have the log-odds formulation of occupancy probability:
@@ A.1 Occupancy Grid Mapping Algorithm
where $v_{i}$ denotes the occupancy probability of $i^{th}$ voxel in the grid $F^{G}_{t}$, $z_{j}$ is the measurement event that $j^{th}$ camera ray passes through this voxel. The numerator of the item $\log\frac{p(z_{j}|v_{i}=1)}{p(z_{j}|v_{i}=0)}$ means that the probability of being passed for a voxel if this voxel is occupied in fact, which shows the confidence of ray casting process. Obviously, the numerator and denominator can be set as an empirical constant. Therefore, we update the log-odds occupancy probability of each voxel in the grid $F^{G}_{t}$ by adding a constant one time when a single camera ray passes through this voxel. Note that the probabilistic occupancy grid $F^{G}$ is continually updated within one episode. Finally, the occupancy status of voxels can be classified into three categories: unknown, occupied, and free, by setting a probability threshold.
We use a constant $C^{\prime}$ to represent the value of a ratio $|\frac{\log p(z_{j}|v_{i}=1)}{\log p(z_{j}|v_{i}=0)}|$ in Eq. 6, where the denominator term $|\log p(z_{j}|v_{i}=0)|$ is set to 0.01. To demonstrate the robustness and optimality of hyper-parameter in our implementation, we show the experimental results in Tab. 5.
Table 5: Evaluation results for different empirical parameters in the implementation of occupancy grid mapping algorithm. Note that we use a constant $C^{\prime}$ to represent the value of a ratio $|\log\frac{p(z_{j}|v_{i}=1)}{p(z_{j}|v_{i}=0)}|$ in Eq. 6
@@ A.2 Network and Training
Training. In Isaac Gym [22], we set 256 parallel environments for policy training, where each environment corresponds with one building-level object. Once the coverage ratio reaches a threshold (99% in practice), a collision happens or the episode length reaches the maximum threshold (100 steps), the environment is reset and the building to be reconstructed is replaced. Our NBV policy is optimized through over 32 million iterations and uses approximately 24 hours of training time on an NVIDIA V100 GPU. All networks are randomly initialized and trained end-to-end.
Network. We propose a multi-source representation that has better generalizability. In particular, we first build two mid-level representations: a 3D geometric representation $F^{G}$ from depth maps and a semantic representation $F^{S}$ from RGB images. Geometric representation $F^{G}$ encoded from depth images is with shape of $(N,X,Y,Z,4)$, where the number of parallel training environments $N$ is 256, grid size $(X,Y,Z)$ is $(20,20,20)$ and last dimension represents the 3D world coordinate and occupancy possibility. Semantic representation $F_{S}$ is stacked grayscale images encoded from multi-view RGB images with the shape of $(N,M,H,W)$, where the number of RGB images $M$ is 5, the size of grayscale images is $(64,64)$
@@
Our policy network PPO, implemented by Stable Baseline3 [29], is a 3-layer multi-layer perceptron (MLP). The output of our policy is used to parameterize a distribution over our 5-dimension action space. In this way, the action can be drawn from the stochastic policy $a\sim\pi(\cdot|o_{t})$.
@@ A.3 Baseline Policies
5) ActiveRMap [48]: The working pipeline is similar to uncertainty-guided policy. In particular, we implement the “discrete (free)” setup of ActiveRMap, which constrains the drone agent on the pre-defined hemisphere. Having considered collision avoidance in the design of viewpoint set, we remove the collision penalty in its optimization objective.
@@
7) Ours with Scan-RL’s Representation: To further compare the former Next-Best-View policy Scan-RL [28] with us, we implement Scan-RL in our experimental setup, with our action and state space. This free-space NBV policy uses Scan-RL’s representation only extracted from RGB images instead of our original hybrid representations.

### v3

@@ 1 Introduction
Recent advances in 3D reconstruction [26, 40, 41] and neural rendering [23, 39, 50] have significantly enhanced the quality of 3D digitization of large scenes, such as buildings and city landmarks [12, 49, 21, 45, 20]. However, image-capturing process still remains time-consuming and labor-intensive. To scan a building-scale scene using a drone, it may take several days’ effort of a professional team to ensure full coverage. Moreover, even professional pilots may miss some areas at the first scan, leading to multiple rounds for rescanning.
To alleviate the effort in manual scanning, active 3D reconstruction algorithms, especially the Next-Best-View (NBV) policy, have emerged as a promising approach to automate view planning. However, typical rule-based NBV policies [17, 25, 48, 28, 10, 30] heavily rely on empirically designed view-selection criteria and hand-crafted action space, such as a hemisphere as shown in Fig. 1, which limit their generalizability to unseen scenes. To enhance generalization capacity, pioneering work on learning-based NBV policy has been proposed [28, 48, 5, 3]. Nevertheless, they are constrained into impractical setups in real-world scenarios, such as object-centric capturing and limited action space for ground robots, resulting in incomplete 3D reconstruction because of significant self-occlusion.
@@
To overcome these two challenges, we propose GenNBV, an end-to-end generalizable NBV policy. It has a novel design of action spaces and state embeddings, which allows free-space exploration and ensures a large coverage when applied to unseen objects. We first extend the former limited action space, like a hemisphere, to 5D free space. As shown in Fig. 1, this new space is composed of a position subspace of approximately $20m$ x $20m$ x $10m$ and an omnidirectional heading subspace. This free space enables our agent drone to scan from any viewpoint. Moreover, with a much larger action space, our RL-based policy can be easily generalized to building-scale geometries without concerning the potential scale shift from training to evaluation. In contrast, many hand-crafted designs [25, 17] limit the available viewpoints to a small predefined set, and thus cannot generalize if there is a drastic change in scales.
Second, in order to generalize a well-trained NBV policy to unseen scenes during evaluation, we propose a novel state representation directly from raw sensory inputs to guide the policy. In contrast to the widely used neural radiance field (NeRF) [25, 48, 17, 36, 30], our representation does not need time-consuming per-scene optimization. Specifically, our NBV policy incorporates a multi-source state embedding, which comprises three key components: a novel geometric embedding, a semantic embedding, and an action embedding. The geometric embedding is derived from multi-view depth maps and represents an encoded probabilistic 3D grid, while the semantic embedding is extracted from RGB images, and the action embedding is an encoded viewpoint sequence. In contrast to 2D representations used in previous RL-based NBV policies [5, 3], our multi-source state embedding, extracted from a probabilistic 3D grid, is a more comprehensive representation that supports a robust prediction of next viewpoints.
To validate the effectiveness of GenNBV, we construct a benchmark for training and evaluation with Houses3K [27] dataset from the NVIDIA Isaac Gym [22] simulator. We further test its generalization ability on novel buildings and object categories from OmniObject3D [44] and Objaverse [7] datasets. Our evaluation metrics quantify the completeness, accuracy, and efficiency of active 3D reconstruction. To consolidate completeness and efficiency into a single number, we also propose a metric, the Area Under the Curve (AUC) of the coverage. Compared to heuristic [13], information gain-based [13, 17, 48], and RL-based baselines [28], our method achieves the best result under different metrics. We also provide visualizations to demonstrate the generalizability of GenNBV. A demo video is attached in the appendix to further illustrate our method.
@@ 3.2 Generalizable State Embedding
To ensure that the policy learned on one set of 3D objects can generalize to objects with different appearances and structures, a smart state embedding of raw sensor observations is needed to capture invariant features across different objects. Previous methods  [28, 3] only extract representations in 2D space, which are very sensitive to appearance changes, and our experiments have shown that policies only trained on 2D features may not be generalized to different objects. Instead, we propose a multi-source representation that has better generalizability.
@@
During each scanning (one episode in RL), at each step $t+1$, we update the probabilistic occupancy grid $F_{t+1}^{G}$ based on the preceding grid $F_{t}^{G}$ and the current observation. Specifically, the grid is updated through Bresenham’s line algorithm [2], which casts the ray path in 3D space between the camera viewpoint and the endpoints among the point cloud back-projected from depth $D_{t+1}$. Following the classical occupancy grid mapping algorithm [42], we have the log-odds formulation of occupancy probability:
@@ A.1 Occupancy Grid Mapping Algorithm
where $v_{i}$ denotes the occupancy probability of $i^{th}$ voxel in the grid $F^{G}_{t}$, $z_{j}$ is the measurement event that $j^{th}$ camera ray passes through this voxel.
For the item $C=\log\frac{p(z_{j}|v_{i}=1)}{p(z_{j}|v_{i}=0)}$, there are only two cases for the measurement event in fact: $z_{j}=0$ or $z_{j}=1$. Thus, if the measurement event $z_{j}$ (i.g., the voxel is passed through by the $j^{th}$ camera ray) happens, we’ll update the occupancy by adding the value of $C_{1}=\log\frac{p(z_{j}=1|v_{i}=1)}{p(z_{j}=1|v_{i}=0)}$. If it’s not passed, we’ll add the value $C_{2}=\log\frac{p(z_{j}=0|v_{i}=1)}{p(z_{j}=0|v_{i}=0)}$. The values of $C_{1}$ and $C_{2}$ can be set as an empirical constant, depending on factors such as the accuracy of ray casting and the confidence of each ray. Actually, $C’=|\frac{C_{1}}{C_{2}}|$. We set a high incremental value for $C^{\prime}$ (i.e., high confidence) because our experiments are based on the realistic simulator and accurate observations like depth maps.
Therefore, we can update the occupancy status of each voxel in the grid $F^{G}_{t}$ by adding a constant for each ray casting process. Note that the probabilistic occupancy grid $F^{G}$ is continuously updated within an episode. Finally, the occupancy status of voxels can be classified into three categories: unknown, occupied, and free, by setting an empirical threshold.
To demonstrate the robustness and optimality of hyper-parameter in our implementation, we show the experimental results in Tab. 5.
Table 5: Evaluation results for different empirical parameters in the implementation of occupancy grid mapping algorithm. Note that we use a constant $C^{\prime}$ to represent the value of a ratio $|\frac{C_{1}}{C_{2}}|$ introduced in A.1
@@ A.2 Network and Training
Training. In Isaac Gym [22], we set 256 parallel environments for policy training, where each environment corresponds with one building-level object. Once the coverage ratio reaches a threshold (99% in practice), a collision occurs, or the episode length reaches the maximum threshold (100 steps), the environment is reset and the building to be reconstructed is replaced. Our NBV policy is optimized through more than 32 million iterations and uses approximately 24 hours of training time on an NVIDIA V100 GPU. All networks are randomly initialized and trained end-to-end.
Network. We propose a multi-source representation that has better generalizability. In particular, we first build two mid-level representations: a 3D geometric representation $F^{G}$ from depth maps and a semantic representation $F^{S}$ from RGB images. The geometric representation $F^{G}$ encoded from depth images is with the shape of $(N,X,Y,Z,4)$, where the number of parallel training environments $N$ is 256, the grid size $(X,Y,Z)$ is $(20,20,20)$, and the last dimension represents the 3D world coordinate and occupancy possibility. Semantic representation $F_{S}$ is stacked grayscale images encoded from multi-view RGB images with the shape of $(N,M,H,W)$, where the number of RGB images $M$ is 5, the size of grayscale images is $(64,64)$
@@
Our policy network PPO, implemented by Stable Baseline3 [29], is a 3-layer multi-layer perceptron (MLP). The output of our policy is used to parameterize a distribution over our 5-dimensional action space. In this way, the action can be drawn from the stochastic policy $a\sim\pi(\cdot|o_{t})$.
@@ A.3 Baseline Policies
5) ActiveRMap [48]: The working pipeline is similar to uncertainty-guided policy. In particular, we implement the “discrete (free)” setup of ActiveRMap, which constrains the drone agent on the pre-defined hemisphere. Having considered collision avoidance in the design of the viewpoint set, we remove the collision penalty in its optimization objective.
@@
7) Ours with Scan-RL’s Representation: To further compare the former Next-Best-View policy Scan-RL [28] with us, we implement Scan-RL in our experimental setup, with our action and state space. This free-space NBV policy uses Scan-RL’s representation, which is only extracted from RGB images instead of our original hybrid representations.